OpenStudy (anonymous):

Judith is digging for sandstone at a geological site. She has 160 meters of rope and 4 stakes to mark off a rectangular area. Which set of dimensions will create a rectangle using all the rope Judith has with her? 16 m × 10 m 80 m × 80 m 65 m × 10 m 70 m × 10 m

OpenStudy (anonymous):

@phi

OpenStudy (phi):

the perimeter (length of the rope when you "wrap" it around the stakes) will be L+L+W+W or 2L+2W or 2(L+W) I would add up the two dimensions (the L and W) of each choice, double it and see if I get 160

OpenStudy (anonymous):

would it be the last answer choice? @phi

OpenStudy (phi):

yes, because 2(70+10) = 2*80= 160

OpenStudy (anonymous):

thank u!